OpenStudy (anonymous):

Solve. 4(a + 3) = 18 – 2(3 – 2a) A. –3 B. –1 C. no solution D. all real numbers

OpenStudy (jadeishere):

Would you like me to walk you through the steps to solving this equation? >.<

OpenStudy (anonymous):

@pinkbubbles

OpenStudy (anonymous):

yuor back

OpenStudy (jadeishere):

Kay 4(a + 3) = 18 – 2(3 – 2a) 4a + 12 = 18 - 6 - 4a 4a + 12 = 4a + 12 all real numbers because any number could be 'a' and it would still equal the other side of the equation :)

OpenStudy (anonymous):

what do you think @pinkbubbles

OpenStudy (anonymous):

i agree